jigglypuff314 (jigglypuff314):

Determine the value of the function \[f(x) = 24x + 24x ^{3/2} - 16x ^{2}\] at the x-coordinate where the second derivative is zero. Well, so far I've got the second derivative to be \[f''(x) = 18x ^{-1/2}-32\] (which I am not even sure is correct) but if it is, how would I find it where "the x-coordinate is zero"?

OpenStudy (anonymous):

So you have: \[f(x)=24x+24x^\frac{3}{2}+16x^2\] So you have the derivative: \[f'(x)=24+36x^\frac{1}{2}+32x\] And the second derivative: \[f''(x)=18x^{-\frac{1}{2}}+32=\frac{18}{\sqrt{x}}+32\] So you want to find the x-value when \([f''(x)=0]\)

OpenStudy (anonymous):

So then you would have: \[\eqalign{ &f''(x)=0 \\ &\frac{18}{\sqrt{x}}+32=0 \\ &\frac{18}{\sqrt{x}}=-32 \\ &\frac{1}{\sqrt{x}}=-\frac{32}{18} \\ &\sqrt{x}=-\frac{18}{32} \\ &x=\left(-\frac{18}{32}\right)^2 \\ }\]

jigglypuff314 (jigglypuff314):

it's supposed to be a -32...

OpenStudy (anonymous):

OOOOOOH OUCH. Sorry yeah so then it would be: \[\eqalign{ &f''(x)=0 \\ &\frac{18}{\sqrt{x}}-32=0 \\ &\frac{18}{\sqrt{x}}=32 \\ &\frac{1}{\sqrt{x}}=\frac{32}{18} \\ &\sqrt{x}=\frac{18}{32} \\ &x=\left(\frac{18}{32}\right)^2 \\ }\]

jigglypuff314 (jigglypuff314):

so x = 0.316?

jigglypuff314 (jigglypuff314):

okay, I got it, thank you! :)

OpenStudy (anonymous):

Ehhh...yeah! About no prob :)